OpenStudy (anonymous):

If a line crosses the y-axis at (0, 1) and has a slope of 45, what is the equation of the line? 4y − 5x = 5 y − 4x = 5 5y + 4x = 5 5y − 4x = 5

OpenStudy (texaschic101):

there is 2 ways to do this. You can either use point-slope formula (y - y1 = m(x - x1), or you can just use slope-intercept formula (y = mx + b) Before I go any further, are you sure your slope is 45 ?

OpenStudy (anonymous):

oh sorry, its 4/5

OpenStudy (texaschic101):

which formula do you want to use....or does it matter ?

OpenStudy (anonymous):

y=mx=b

OpenStudy (anonymous):

y=mx+b

OpenStudy (texaschic101):

y = mx + b slope(m) = 4/5 (0,1)...x = 0 and y = 1 now we sub 1 = 4/5(0) + b 1 = b y = 4/5x + 1 --- multiply everything by 5 to get rid of the fraction 5y = 4x + 5 -- subtract 4x from both sides 5y - 4x = 5 <== your equation

OpenStudy (anonymous):

Thank you!